# OE_^ OmOE_^ OmOE_^ Omm!OE_^ mmmmOE_^

### v1

@@ 6.1. Hyperbolic porosity
For $u=(u_{1},\ldots,u_{n})\in\mathbb{R}^{n}$, define the vector field
$U^{\pm}u\coloneqq u_{1}U_{1}^{\pm}+\cdots+u_{n}U_{n}^{\pm}.$
@@
We say that $\Omega\subset S^{*}M$ is hyperbolic (±$(\pm\nu,\mp\varepsilon)$-porous on balls from scales $\alpha_{0}$ to $\alpha_{1}$ if for all $(x,\xi)\in S^{*}M$, if $q\in\pi_{S}^{-1}((x,\xi))$ and $\alpha\in[\alpha_{0},\alpha_{1}]$, there exists $u_{0}\in\mathbb{R}^{n}$ with $|u_{0}|\leq\alpha$ such that
$\pi_{S}\{e^{\mathcal{V}^{\mp}v}e^{U^{\pm}(u+u_{0})}q:u\in\mathbb{R}^{n},|u|\leq\nu\alpha,v\in\mathbb{R}^{n+1},|v|\leq\varepsilon\}\cap\Omega=\emptyset.$
@@
We say that $\Omega\subset S^{*}M$ is hyperbolic (±$(\pm\nu,\mp\varepsilon)$-porous on lines from scales $\alpha_{0}$ to $\alpha_{1}$ if for all $(x,\xi)\in S^{*}M$, if $q\in\pi_{S}^{-1}((x,\xi))$ and $\alpha\in[\alpha_{0},\alpha_{1}]$, there exists $t_{0}\in[-\alpha,\alpha]$ such that
$\pi_{S}\{(e^{\mathcal{V}^{\mp}v}e^{U^{\pm}u+U_{1}^{\pm}t_{0}}q):u\in\mathbb{R}^{n},|u|\leq\nu\alpha,v\in\mathbb{R}^{n+1},|v|\leq\varepsilon\}\cap\Omega=\emptyset.$
@@
Figure 1. The left image represents the set from Definition 6.1: $\{e^{\mathcal{V}^{-}v}e^{U^{+}u}q_{0}:u\in\mathbb{R}^{n},|u|\leq\nu\alpha,v\in\mathbb{R}^{n+1},|v|\leq\varepsilon\}$, where $q_{0}=e^{U^{+}u_{0}}q$. The right image represents the set from Definition 6.2: $\{e^{\mathcal{V}^{-}v}e^{U^{+}u}q_{1}:u\in\mathbb{R}^{n},|u|\leq\nu\alpha,v\in\mathbb{R}^{n+1},|v|\leq\varepsilon\}$, where $q_{1}=e^{U^{+}_{1}t_{0}}q$. The picture is to give intuition only; as $U^{+}_{i}$, $U_{i}^{-}$, $X$ do not commute, they do not give a coordinate system.
@@
$a_{+}=b\prod_{k=1}^{T_{1}}\left(\chi_{w_{k}^{+}}\circ\varphi_{-k}\right)\quad\text{and}\quad a_{-}=b\prod_{k=0}^{T_{1}-1}\left(\chi_{w_{k}^{-}}\circ\varphi_{k}\right),$
@@
Clearly, $\operatorname{supp}a_{\pm}\cap S^{*}M=\mathcal{A}_{\pm}\cap S^{*}M$ and $\mathcal{A}_{\pm}$ are invariant under scalings of $\xi$ within the compact set $\{\tfrac{1}{4}\leq|\xi|_{g}\leq 4\}$. Later, we exploit this homogeneity of $\mathcal{A}_{\pm}$ and replace $\operatorname{supp}a_{\pm}$ with $\mathcal{A}_{\pm}$.
@@ 6.1.1. Ball porosity
For each open set $U\subset S^{*}M$, there exists some $T>0$ such that for all $q\in F^{*}M$ and $t\geq T$, we have
$\varphi_{-t}\pi_{S}\{e^{U^{+}u}q:u\in\mathbb{R}^{n},|u|\leq 1\}\cap U\neq\emptyset.$
@@
$\operatorname{supp}f\subset\pi_{S}\left\{e^{\mathcal{V}^{-}v}e^{U^{+}u}q_{0}:u\in\mathbb{R}^{n},|u|\leq 1,v\in\mathbb{R}^{n+1},|v|\leq\varepsilon\right\}.$
@@
$\varphi_{-t}(\operatorname{supp}f)$ $\subset\varphi_{-t}\pi_{S}\left\{e^{\mathcal{V}^{-}v}e^{U^{+}u}q_{0}:u\in\mathbb{R}^{n},|u|\leq 1,v\in\mathbb{R}^{n+1},|v|\leq\varepsilon\right\}$
$=\pi_{S}\left\{\varphi_{-t}(e^{\mathcal{V}^{-}v}e^{U^{+}u}q_{0}):u\in\mathbb{R}^{n},|u|\leq 1,v\in\mathbb{R}^{n+1},|v|\leq\varepsilon\right\}$
$\subset\pi_{S}\left\{e^{\mathcal{V}^{-}v}e^{U^{+}u}\varphi_{-t}(q_{0}):u\in\mathbb{R}^{n},|u|\leq e^{t},v\in\mathbb{R}^{n+1},|v|\leq\varepsilon\right\}.$
Therefore, for some $|u_{0}|\leq e^{t}$ and $|v_{0}|\leq\varepsilon$, $e^{\mathcal{V}^{-}v_{0}}e^{U^{+}u_{0}}\varphi_{-t}(q_{0})\in\pi_{S}^{-1}\operatorname{supp}\chi$. By the choice of $\varepsilon$, $\pi_{S}(e^{U^{+}u_{0}}\varphi_{-t}(q_{0}))\in U$.
From (6.4), $\pi_{S}(e^{U^{+}u_{0}}\varphi_{-t}(q_{0}))=\varphi_{-t}(\pi_{S}(e^{U^{+}e^{-t}u_{0}}q_{0}))$. Clearly, $|e^{-t}u_{0}|\leq 1$.
@@
We now prove the hyperbolic ball porosity of $\operatorname{supp}a_{+}\cap S^{*}M$. The following proof takes inspiration from  [DJ18]*Lemma 5.10.
@@
There exists $\nu_{1},\varepsilon_{0},K_{1}>0$, where $\nu_{1}=e^{-T-1}\varepsilon_{0}$ for some $T>1$, depending only on $M,a_{1},a_{2}$, such that $\operatorname{supp}a_{+}\cap S^{*}M$ is hyperbolic $(+\nu_{1},-\varepsilon_{0})$-porous on balls from scales $K_{1}h^{\rho}$ to $1$ in the sense of Definition  6.1.
@@
From (5.3), there exists $U_{1},U_{2}\subset S^{*}M$ such that for $w=1,2$, $U_{w}\Subset S^{*}M\setminus\operatorname{supp}a_{w}$.
@@
(6.5) $\pi_{S}(\varphi_{-T}(e^{U^{+}u_{w}}(q)))\in U_{w}.$
Set $K_{1}\coloneqq e^{2+T}$. Fix $q_{0}\in F^{*}M$ and $\alpha\in[K_{1}h^{\rho},1]$. Let $t$ be the unique integer such that $e^{-t}<\alpha\leq e^{-t+1}$. From  (5.7), $11<T+t\leq T_{1}-$ 1. Set $w=w_{t+T}^{+}\in\{1,2\}$ and note that $\operatorname{supp}a_{+}\cap\varphi_{t+T}(S^{*}M\setminus\operatorname{supp}a_{w})=\emptyset$.
@@
(6.6) $\pi_{S}(\varphi_{-T-t}(e^{U^{+}e^{-t}u_{0}}q_{0}))=\pi_{S}(\varphi_{-T}(e^{U^{+}u_{0}}\varphi_{-t}(q_{0})))\in U_{w}.$
@@
$\pi_{S}(e^{\mathcal{V}^{-}v}e^{U^{+}u}q)\subset S^{*}M\setminus\operatorname{supp}a_{w}.$
@@
(6.7) $\{e^{\mathcal{V}^{-}v}e^{U^{+}u}(\varphi_{-T-t}(e^{U^{+}e^{-t}u_{0}}q_{0})):|u|,|v|\leq\varepsilon_{0}\}\subset\pi_{S}^{-1}(S^{*}M\setminus\operatorname{supp}a_{w}).$
@@
(6.8) $e^{\mathcal{V}^{-}v}e^{U^{+}u}\varphi_{t}(q)=\varphi_{t}(e^{\mathcal{V}^{-}v^{\prime}}e^{U^{+}e^{t}u}q),$
@@
$\pi_{S}\left\{\varphi_{-t-T}(e^{\mathcal{V}^{-}v}e^{U^{+}(e^{-T-t}u+e^{-t}u_{0})}q_{0}):|u|,|v|\leq\varepsilon_{0}\right\}\subset S^{*}M\setminus\operatorname{supp}a_{w}.$
@@
$\pi_{S}\left\{e^{\mathcal{V}^{-}v}e^{U^{+}(u+e^{-t}u_{0})}q_{0}:|u|\leq\nu_{1}\alpha,|v|\leq\varepsilon_{0}\right\}\subset\varphi_{t+T}\left(S^{*}M\setminus\operatorname{supp}a_{w}\right)\subset S^{*}M\setminus\operatorname{supp}a_{+},$
@@ 6.1.2. Line porosity
We show that $\operatorname{supp}a_{-}\cap S^{*}M$ is hyperbolic porous on lines. This result uses our assumption that $U_{\mu}$ is $U_{1}^{-}$ dense. Similarly to Lemma 6.6, the proof is adapted from  [DJ18]*Lemma 5.10.
@@
There exists $\nu_{1},\varepsilon_{0},K_{1}>0$, with $\nu_{1}=\varepsilon_{0}/(3T)$ for some $T>1$, depending only on $M,a_{1},a_{2}$ such that $\operatorname{supp}a_{-}\cap S^{*}M$ is hyperbolic $(-\nu_{1},+\varepsilon_{0})$-porous on lines from scales $K_{1}h^{\rho}$ to $1$ in the sense of Definition  6.2.
@@
For $w=1,2$, from (5.3), there exists a $U_{1}^{-}$-dense set $U_{w}\subset S^{*}M$ such that $U_{w}\Subset S^{*}M\setminus\operatorname{supp}a_{w}$.
@@
$\pi_{S}(e^{\mathcal{V}^{+}v}e^{U^{-}u}q)\subset S^{*}M\setminus\operatorname{supp}a_{w}.$
@@
$\pi_{S}\left\{e^{\mathcal{V}^{+}v}e^{U^{-}u}\varphi_{j}(e^{t_{0}U_{1}^{-}}q_{0}):|u|,|v|\leq\varepsilon_{0}\right\}\subset S^{*}M\setminus\operatorname{supp}a_{w}.$
@@
$e^{\mathcal{V}^{+}v}e^{U^{-}u}\varphi_{j}(q)=\varphi_{j}(e^{\mathcal{V}^{+}v^{\prime}}e^{U^{-}e^{-j}u}q),$
@@
$\pi_{S}\{e^{\mathcal{V}^{+}v}e^{U^{-}u}e^{t_{0}U_{1}^{-}}q_{0}:|u|\leq\nu_{1}\alpha,|v|<\varepsilon_{0}\}\subset\varphi_{-j}(S^{*}M\setminus\operatorname{supp}a_{w})\subset S^{*}M\setminus\operatorname{supp}a_{-},$
@@ 6.2. Symplectomorphisms on hyperbolic space
The use of two symplectomorphisms is necessary; as symplectomorphisms are volume preserving, we cannot use a single diffeomorphism to simultaneously straighten out both $L_{s}$ and $L_{u}$.
@@
Denote by $\mathcal{P}(x,y)$ the $2$-dimensional version of the Poisson kernel, defined on the ball model of $\mathbb{H}^{n+1}$ by
@@ 6.3. Hyperbolic fractal uncertainty principle
The goal of this subsection is to prove Proposition 6.10, stated following this paragraph. Intuitively, this proposition tell us that the hyperbolic $(+\nu_{1},-\varepsilon_{0})$-porosity on balls of $\operatorname{supp}a_{+}$ from Lemma 6.6 and the hyperbolic $(-\nu_{1},+\varepsilon_{0})$-porosity on lines of $\operatorname{supp}a_{-}$ from Lemma 6.7 give a fractal uncertainty principle. The statement and the proof of Proposition 6.10 are adapted to higher dimensions and to manifolds from  [DJ18]*Proposition 5.7. Note that Lemma 6.6 and Lemma 6.7 give two sets of values for $\nu_{1}$, $\varepsilon_{0}$, and $K_{1}$. From now on, we assume that $\nu_{1}$ and $\varepsilon_{0}$ are each equal to the minimum of their two values and $K_{1}$ is equal to the maximum of its two values. We can write $\varepsilon_{0}=C\nu_{1}$, where $C$ depends only on $M,a_{1},a_{2}$.
@@
Recall the definition of $B_{-}(x,\xi)$ from  (6.11). Then since $U_{i}^{\pm}$ generate the horocyclic flows,
@@
By Lemma 6.6, we know $\operatorname{supp}a_{+}\cap S^{*}M$ is hyperbolic $(+\nu_{1},-\varepsilon_{0})$-porous on balls from scales $K_{1}h^{\rho}$ to $1$. Thus, there exists $u_{0}\in\mathbb{R}^{n}$, $|u_{0}|\leq\alpha$ such that
$\{\pi_{S}(e^{\mathcal{V}^{-}v}e^{U^{+}(u+u_{0})}(x_{0},\xi_{0},\Xi_{0})):|u|\leq\nu_{1}\alpha,|v|\leq\varepsilon_{0}\}\cap\operatorname{supp}a_{+}=\emptyset.$
As $(\mathcal{A}_{+}\cap S^{*}M)=(\operatorname{supp}a_{+}\cap S^{*}M)$ and $\mathcal{A}_{+}$ is homogeneous inside $\{\tfrac{1}{4}\leq|\xi|_{g}\leq 4\}$,
(6.20) $\{f_{\lambda}(\pi_{S}(e^{\mathcal{V}^{-}v}e^{U^{+}(u+u_{0})}(x_{0},\xi_{0},\Xi_{0}))):|u|\leq\nu_{1}\alpha,|v|\leq\varepsilon_{0},\tfrac{1}{4}\leq\lambda\leq 4\}\cap\mathcal{A}_{+}=\emptyset,$
@@
$(x_{1},\xi_{1},\Xi_{1})\coloneqq e^{U^{+}u_{0}}(x_{0},\xi_{0},\Xi_{0}),\quad(x_{1},\xi_{1})\in V^{\prime\prime}.$
@@
$\Theta:\tilde{U}\times\left[-\tfrac{1}{4}-\tfrac{2\nu_{1}}{C_{1}^{2}},4+\tfrac{2\nu_{1}}{C_{1}^{2}}\right]\rightarrow W^{\prime\prime},\quad(u,v,\lambda)\mapsto\kappa_{0}^{+}(f_{\lambda}(\pi_{S}(e^{\mathcal{V}^{-}v}e^{U^{+}u}(x_{0},\xi_{0},\Xi_{0})))),$
@@
We now show that the hyperbolic $(-\nu_{1},+\varepsilon_{0})$-porosity on lines of $\operatorname{supp}a_{-}$ from Lemma 6.7 implies that $\tilde{\mathcal{A}}_{-}$ has a property similar to line porosity in Definition 4.2. Similarly to Lemma 6.11, the following lemma is adapted from  [DJ18]*Lemma 5.8.
@@
By Lemma 6.7, $\operatorname{supp}a_{-}\cap S^{*}M$ is hyperbolic $(-\nu_{1},+\varepsilon_{0})$-porous on lines from scales $K_{1}h^{\rho}$ to $1$. Thus, there exists $|t_{0}|\leq\alpha$ such that
$\left\{\pi_{S}(e^{\mathcal{V}^{+}v}e^{U^{-}u+t_{0}U_{1}^{-}}(x(0),\xi(0),\Xi_{0})):|u|\leq\nu_{1}\alpha,|v|\leq\varepsilon_{0}\right\}\cap\operatorname{supp}a_{-}=\emptyset.$
As $(\mathcal{A}_{-}\cap S^{*}M)=(\operatorname{supp}a_{-}\cap S^{*}M)$ and $\mathcal{A}_{-}$ is homogeneous inside $\{\tfrac{1}{4}\leq|\xi|_{g}\leq 4\}$,
(6.25) $\left\{f_{\lambda}(\pi_{S}(e^{\mathcal{V}^{+}v}e^{U^{-}u+t_{0}U_{1}^{-}}(x(0),\xi(0),\Xi_{0}))):|u|\leq\nu_{1}\alpha,|v|\leq\varepsilon_{0},\tfrac{1}{4}\leq\lambda\leq 4\right\}\cap\mathcal{A}_{-}=\emptyset,$

### v2

@@ 6.1. Hyperbolic porosity
For $u=(u_{1},\ldots,u_{n})\in\mathbb{R}^{n}$, define the vector field on $F^{*}M$
$\mathcal{U}^{\pm}u\coloneqq u_{1}U_{1}^{\pm}+\cdots+u_{n}U_{n}^{\pm}.$
@@
We say that $\Omega\subset S^{*}M$ is hyperbolic $(\nu^{\pm},\varepsilon^{\mp})$-porous on balls from scales $\alpha_{0}$ to $\alpha_{1}$ if for all $q\in F^{*}M$ and $\alpha\in[\alpha_{0},\alpha_{1}]$, there exists $u_{0}\in\mathbb{R}^{n}$ with $|u_{0}|\leq\alpha$ such that
$\pi_{S}\{e^{\mathcal{V}^{\mp}v}e^{\mathcal{U}^{\pm}(u+u_{0})}q:u\in\mathbb{R}^{n},|u|\leq\nu\alpha,v\in\mathbb{R}^{n+1},|v|\leq\varepsilon\}\cap\Omega=\emptyset.$
@@
We say that $\Omega\subset S^{*}M$ is hyperbolic $(\nu^{\pm},\varepsilon^{\mp})$-porous on lines from scales $\alpha_{0}$ to $\alpha_{1}$ if for all $q\in F^{*}M$ and $\alpha\in[\alpha_{0},\alpha_{1}]$, there exists $t_{0}\in[-\alpha,\alpha]$ such that
$\pi_{S}\{(e^{\mathcal{V}^{\mp}v}e^{\mathcal{U}^{\pm}u+U_{1}^{\pm}t_{0}}q):u\in\mathbb{R}^{n},|u|\leq\nu\alpha,v\in\mathbb{R}^{n+1},|v|\leq\varepsilon\}\cap\Omega=\emptyset.$
@@
Figure 1. The left image represents the set from Definition 6.1: $\{e^{\mathcal{V}^{-}v}e^{\mathcal{U}^{+}u}q_{0}:u\in\mathbb{R}^{n},|u|\leq\nu\alpha,v\in\mathbb{R}^{n+1},|v|\leq\varepsilon\}$, where $q_{0}=e^{\mathcal{U}^{+}u_{0}}q$. The right image represents the set from Definition 6.2: $\{e^{\mathcal{V}^{-}v}e^{\mathcal{U}^{+}u}q_{1}:u\in\mathbb{R}^{n},|u|\leq\nu\alpha,v\in\mathbb{R}^{n+1},|v|\leq\varepsilon\}$, where $q_{1}=e^{U^{+}_{1}t_{0}}q$. The picture is to give intuition only; as $U^{+}_{i}$, $U_{i}^{-}$, $X$ do not commute, they do not give a coordinate system.
@@
$a_{+}=b^{T_{1}-1}\prod_{k=1}^{T_{1}}\left(\chi_{w_{k}^{+}}\circ\varphi_{-k}\right)\quad\text{and}\quad a_{-}=b^{T_{1}-1}\prod_{k=0}^{T_{1}-1}\left(\chi_{w_{k}^{-}}\circ\varphi_{k}\right),$
@@
Clearly, $\mathcal{A}_{\pm}$ are invariant under scalings of $\xi$ within the compact set $\{\tfrac{1}{4}\leq|\xi|_{g}\leq 4\}$. We will use this fact and use $\mathcal{A}_{\pm}$ in lieu of $\operatorname{supp}a_{\pm}$ in our argument.
@@ 6.1.1. Ball porosity
For each open nonempty set $U\subset S^{*}M$, there exists some $T>0$ such that for all $q\in F^{*}M$ and $t\geq T$, we have
$\varphi_{-t}\pi_{S}\{e^{\mathcal{U}^{+}u}q:u\in\mathbb{R}^{n},|u|\leq 1\}\cap U\neq\emptyset.$
@@
$\operatorname{supp}f\subset\pi_{S}\left\{e^{\mathcal{V}^{-}v}e^{\mathcal{U}^{+}u}q_{0}:u\in\mathbb{R}^{n},|u|\leq 1,v\in\mathbb{R}^{n+1},|v|\leq\varepsilon\right\}.$
@@
$\varphi_{-t}(\operatorname{supp}f)$ $\subset\varphi_{-t}\pi_{S}\left\{e^{\mathcal{V}^{-}v}e^{\mathcal{U}^{+}u}q_{0}:u\in\mathbb{R}^{n},|u|\leq 1,v\in\mathbb{R}^{n+1},|v|\leq\varepsilon\right\}$
$=\pi_{S}\left\{\varphi_{-t}(e^{\mathcal{V}^{-}v}e^{\mathcal{U}^{+}u}q_{0}):u\in\mathbb{R}^{n},|u|\leq 1,v\in\mathbb{R}^{n+1},|v|\leq\varepsilon\right\}$
$\subset\pi_{S}\left\{e^{\mathcal{V}^{-}v}e^{\mathcal{U}^{+}u}\varphi_{-t}(q_{0}):u\in\mathbb{R}^{n},|u|\leq e^{t},v\in\mathbb{R}^{n+1},|v|\leq\varepsilon\right\}.$
Therefore, for some $|u_{0}|\leq e^{t}$ and $|v_{0}|\leq\varepsilon$, $e^{\mathcal{V}^{-}v_{0}}e^{\mathcal{U}^{+}u_{0}}\varphi_{-t}(q_{0})\in\pi_{S}^{-1}\operatorname{supp}\chi$. By the choice of $\varepsilon$, $\pi_{S}(e^{\mathcal{U}^{+}u_{0}}\varphi_{-t}(q_{0}))\in U$.
From (6.4), $\pi_{S}(e^{\mathcal{U}^{+}u_{0}}\varphi_{-t}(q_{0}))=\varphi_{-t}(\pi_{S}(e^{\mathcal{U}^{+}e^{-t}u_{0}}q_{0}))$. Clearly, $|e^{-t}u_{0}|\leq 1$.
@@
We now prove the hyperbolic ball porosity of $\mathcal{A}_{+}\cap S^{*}M$. The following proof takes inspiration from  [DJ18]*Lemma 5.10.
@@
There exists $\nu_{1},\varepsilon_{0},K_{1}>0$, where $\nu_{1}=e^{-T-1}\varepsilon_{0}$ for some $T>1$, depending only on $M,a_{1},a_{2}$, such that $\mathcal{A}_{+}\cap S^{*}M$ is hyperbolic $(\nu_{1}^{+},\varepsilon_{0}^{-})$-porous on balls from scales $K_{1}h^{\rho}$ to $1$ in the sense of Definition  6.1.
@@
From (5.3), there exist $U_{1},U_{2}\subset S^{*}M$ such that for $w=1,2$, $U_{w}\Subset S^{*}M\setminus\operatorname{supp}a_{w}$. Therefore, $U_{w}\Subset S^{*}M\setminus\operatorname{supp}\chi_{w}$.
@@
(6.5) $\pi_{S}(\varphi_{-T}(e^{\mathcal{U}^{+}u_{w}}(q)))\in U_{w}.$
Set $K_{1}\coloneqq e^{2+T}$. Fix $q_{0}\in F^{*}M$ and $\alpha\in[K_{1}h^{\rho},1]$. Let $t$ be the unique integer such that $e^{-t}<\alpha\leq e^{-t+1}$. From  (5.7), $11<T+t\leq T_{1}-$ 1. Set $w=w_{t+T}^{+}\in\{1,2\}$ and note that $\mathcal{A}_{+}\cap\varphi_{t+T}(S^{*}M\setminus\operatorname{supp}\chi_{w})=\emptyset$.
@@
(6.6) $\pi_{S}(\varphi_{-T-t}(e^{\mathcal{U}^{+}e^{-t}u_{0}}q_{0}))=\pi_{S}(\varphi_{-T}(e^{\mathcal{U}^{+}u_{0}}\varphi_{-t}(q_{0})))\in U_{w}.$
@@
$\pi_{S}(e^{\mathcal{V}^{-}v}e^{\mathcal{U}^{+}u}q)\subset S^{*}M\setminus\operatorname{supp}\chi_{w}.$
@@
(6.7) $\{e^{\mathcal{V}^{-}v}e^{\mathcal{U}^{+}u}(\varphi_{-T-t}(e^{\mathcal{U}^{+}e^{-t}u_{0}}q_{0})):|u|,|v|\leq\varepsilon_{0}\}\subset\pi_{S}^{-1}(S^{*}M\setminus\operatorname{supp}\chi_{w}).$
@@
(6.8) $e^{\mathcal{V}^{-}v}e^{\mathcal{U}^{+}u}\varphi_{t}(q)=\varphi_{t}(e^{\mathcal{V}^{-}v^{\prime}}e^{\mathcal{U}^{+}e^{t}u}q),$
@@
$\pi_{S}\left\{\varphi_{-t-T}(e^{\mathcal{V}^{-}v}e^{\mathcal{U}^{+}(e^{-T-t}u+e^{-t}u_{0})}q_{0}):|u|,|v|\leq\varepsilon_{0}\right\}\subset S^{*}M\setminus\operatorname{supp}\chi_{w}.$
@@
$\pi_{S}\left\{e^{\mathcal{V}^{-}v}e^{\mathcal{U}^{+}(u+e^{-t}u_{0})}q_{0}:|u|\leq\nu_{1}\alpha,|v|\leq\varepsilon_{0}\right\}\subset\varphi_{t+T}\left(S^{*}M\setminus\operatorname{supp}a_{w}\right)\subset S^{*}M\setminus\mathcal{A}_{+},$
@@ 6.1.2. Line porosity
We show that $\mathcal{A}_{-}\cap S^{*}M$ is hyperbolic porous on lines. This result uses our assumption that $U_{\mu}$ is $U_{1}^{-}$ dense. Similarly to Lemma 6.6, the proof is adapted from  [DJ18]*Lemma 5.10.
@@
There exists $\nu_{1},\varepsilon_{0},K_{1}>0$, with $\nu_{1}=\varepsilon_{0}/T$ for some $T>1$, depending only on $M,a_{1},a_{2}$ such that $\mathcal{A}_{-}\cap S^{*}M$ is hyperbolic $(\nu_{1}^{-},\varepsilon_{0}^{+})$-porous on lines from scales $K_{1}h^{\rho}$ to $1$ in the sense of Definition  6.2.
@@
For $w=1,2$, from (5.3), there exists a $U_{1}^{-}$-dense set $U_{w}\subset S^{*}M$ such that $U_{w}\Subset S^{*}M\setminus\operatorname{supp}a_{w}$. Therefore, $U_{w}\Subset S^{*}M\setminus\operatorname{supp}\chi_{w}$.
@@
$\pi_{S}(e^{\mathcal{V}^{+}v}e^{\mathcal{U}^{-}u}q)\subset S^{*}M\setminus\operatorname{supp}\chi_{w}.$
@@
$\pi_{S}\left\{e^{\mathcal{V}^{+}v}e^{\mathcal{U}^{-}u}\varphi_{j}(e^{t_{0}U_{1}^{-}}q_{0}):|u|,|v|\leq\varepsilon_{0}\right\}\subset S^{*}M\setminus\operatorname{supp}\chi_{w}.$
@@
$e^{\mathcal{V}^{+}v}e^{\mathcal{U}^{-}u}\varphi_{j}(q)=\varphi_{j}(e^{\mathcal{V}^{+}v^{\prime}}e^{\mathcal{U}^{-}e^{-j}u}q),$
@@
$\pi_{S}\{e^{\mathcal{V}^{+}v}e^{\mathcal{U}^{-}u}e^{t_{0}U_{1}^{-}}q_{0}:|u|\leq\nu_{1}\alpha,|v|<\varepsilon_{0}\}\subset\varphi_{-j}(S^{*}M\setminus\operatorname{supp}a_{w})\subset S^{*}M\setminus\mathcal{A}_{-},$
@@ 6.2. Symplectomorphisms on hyperbolic space
The use of two symplectomorphisms is necessary; we cannot use a single diffeomorphism to simultaneously straighten out both $L_{s}$ and $L_{u}$.
@@
Denote by $\mathcal{P}(x,y)$ the following function defined on the ball model of $\mathbb{H}^{n+1}$ by
@@ 6.3. Hyperbolic fractal uncertainty principle
The goal of this subsection is to prove Proposition 6.10, stated following this paragraph. Intuitively, this proposition tell us that the hyperbolic $(\nu_{1}^{+},\varepsilon_{0}^{-})$-porosity on balls of $\mathcal{A}_{+}$ from Lemma 6.6 and the hyperbolic $(\nu_{1}^{-},\varepsilon_{0}^{+})$-porosity on lines of $\mathcal{A}_{-}$ from Lemma 6.7 give a fractal uncertainty principle. The statement and the proof of Proposition 6.10 are adapted to higher dimensions and to manifolds from  [DJ18]*Proposition 5.7. Note that Lemma 6.6 and Lemma 6.7 give two sets of values for $\nu_{1}$, $\varepsilon_{0}$, and $K_{1}$. From now on, we assume that $\nu_{1}$ and $\varepsilon_{0}$ are each equal to the minimum of their two values and $K_{1}$ is equal to the maximum of its two values. We can write $\varepsilon_{0}=C\nu_{1}$, where $C$ depends only on $M,a_{1},a_{2}$.
@@
Recall the definition of $B_{-}(x,\xi)$ from  (6.11). Then since $U_{i}^{\pm}$ generate one-parameter unipotent flows,
@@
By Lemma 6.6, we know $\mathcal{A}_{+}\cap S^{*}M$ is hyperbolic $(\nu_{1}^{+},\varepsilon_{0}^{-})$-porous on balls from scales $K_{1}h^{\rho}$ to $1$. Thus, there exists $u_{0}\in\mathbb{R}^{n}$, $|u_{0}|\leq\alpha$ such that
$\{\pi_{S}(e^{\mathcal{V}^{-}v}e^{\mathcal{U}^{+}(u+u_{0})}(x_{0},\xi_{0},\Xi_{0})):|u|\leq\nu_{1}\alpha,|v|\leq\varepsilon_{0}\}\cap\mathcal{A}_{+}=\emptyset.$
As $\mathcal{A}_{+}$ is homogeneous inside $\{\tfrac{1}{4}\leq|\xi|_{g}\leq 4\}$,
(6.20) $\{f_{\lambda}(\pi_{S}(e^{\mathcal{V}^{-}v}e^{\mathcal{U}^{+}(u+u_{0})}(x_{0},\xi_{0},\Xi_{0}))):|u|\leq\nu_{1}\alpha,|v|\leq\varepsilon_{0},\tfrac{1}{4}\leq\lambda\leq 4\}\cap\mathcal{A}_{+}=\emptyset,$
@@
$(x_{1},\xi_{1},\Xi_{1})\coloneqq e^{\mathcal{U}^{+}u_{0}}(x_{0},\xi_{0},\Xi_{0}),\quad(x_{1},\xi_{1})\in V^{\prime\prime}.$
@@
$\Theta:\tilde{U}\times\left[-\tfrac{1}{4}-\tfrac{2\nu_{1}}{C_{1}^{2}},4+\tfrac{2\nu_{1}}{C_{1}^{2}}\right]\rightarrow W^{\prime\prime},\quad(u,v,\lambda)\mapsto\kappa_{0}^{+}(f_{\lambda}(\pi_{S}(e^{\mathcal{V}^{-}v}e^{\mathcal{U}^{+}u}(x_{0},\xi_{0},\Xi_{0})))),$
@@
We now show that the hyperbolic $(\nu_{1}^{-},\varepsilon_{0}^{+})$-porosity on lines of $\mathcal{A}_{-}$ from Lemma 6.7 implies that $\tilde{\mathcal{A}}_{-}$ has a property similar to line porosity in Definition 4.2. Similarly to Lemma 6.11, the following lemma is adapted from  [DJ18]*Lemma 5.8.
@@
By Lemma 6.7, $\mathcal{A}_{-}\cap S^{*}M$ is hyperbolic $(\nu_{1}^{-},\varepsilon_{0}^{+})$-porous on lines from scales $K_{1}h^{\rho}$ to $1$. Thus, there exists $|t_{0}|\leq\alpha$ such that
$\left\{\pi_{S}(e^{\mathcal{V}^{+}v}e^{\mathcal{U}^{-}u+t_{0}U_{1}^{-}}(x(0),\xi(0),\Xi_{0})):|u|\leq\nu_{1}\alpha,|v|\leq\varepsilon_{0}\right\}\cap\mathcal{A}_{-}=\emptyset.$
As $\mathcal{A}_{-}$ is homogeneous inside $\{\tfrac{1}{4}\leq|\xi|_{g}\leq 4\}$,
(6.25) $\left\{f_{\lambda}(\pi_{S}(e^{\mathcal{V}^{+}v}e^{\mathcal{U}^{-}u+t_{0}U_{1}^{-}}(x(0),\xi(0),\Xi_{0}))):|u|\leq\nu_{1}\alpha,|v|\leq\varepsilon_{0},\tfrac{1}{4}\leq\lambda\leq 4\right\}\cap\mathcal{A}_{-}=\emptyset,$